OpenStudy (anonymous):

5-(6k+1)=2[(5k-3)-(k-2)]

OpenStudy (anonymous):

I am wondering why you keep posting the same type of problem. What problem are u having with these?

OpenStudy (anonymous):

there confusing to me I had them last semester and I had the same prob I need more in depth study with them because im just not getting them the smaller ones r simple but the ones that have 2 equal signs and parthesis throw me off.

OpenStudy (anonymous):

OK, I can see how the brackets might cause a some confusion. Can you give me an example of a problem with 2 equals signs in it?

OpenStudy (anonymous):

Why not use www.wolframalpha.com It will break the problem out for you so you can understand it.

OpenStudy (anonymous):

Not equal signs I mean fractions and with negative's. sorry I got confused

OpenStudy (anonymous):

thank u gamma

OpenStudy (anonymous):

I just put this problem into Wolfram, the answer doesn't seem very helpful for an understanding.

OpenStudy (anonymous):

Is the problem in what order to do the simplification that causes the problem to begin with?

OpenStudy (anonymous):

Or something else?

OpenStudy (anonymous):

What I am asking is What do you do first when you see one of these problems?

OpenStudy (anonymous):

I try to break down varables I dont know much how to solve big equations like that so I just dont know I took pre algebra in a 6 week course I passed thank God this is algebra in a 6 week course so theres really not much teaching its fast paced I need to hire a tuttor

OpenStudy (anonymous):

A good first thing to do is to look at "nested" brackets (that means brackets inside brackets) and see if you can't make what is inside them simpler.

OpenStudy (anonymous):

In this case, that's this piece (5k-3)-(k-2)

OpenStudy (anonymous):

Can u do that part?

OpenStudy (anonymous):

ok im writing this down so I understand

OpenStudy (anonymous):

Take your time...

OpenStudy (anonymous):

I guess you have to have some understanding of math to understand what wolfram is doing... Because it does simplify it for you... then its simple to solve. But alas... Oh well.

OpenStudy (anonymous):

@Gammagooo I do agree Wolfram is very useful, especially for checking your work. As you say, need to get past first base and then...

OpenStudy (anonymous):

ok compare variables?

OpenStudy (anonymous):

estudier are you a mommy?

OpenStudy (anonymous):

I would like to know if you can do this part (5k-3)-(k-2)

OpenStudy (anonymous):

k-5?

OpenStudy (anonymous):

No, I was a teacher once...

OpenStudy (anonymous):

I can tell.

OpenStudy (anonymous):

is it k-5?

OpenStudy (anonymous):

Wolfram alpha is best avoided if you don't know how to solve the problems yourself. It is great for checking work (and occasionally getting a foothold on problems), but if you use it to avoid developing an understanding of the problems and the skills it can destroy you later, sadly. Anyways! The parenthesis are best dealt with in these types of problems by expanding first. What you need to do if there's a constant in front of a parenthesis is multiply this constant by each element within the parenthesis, for example, 2[(5k - 3) - (k - 2)] would become... 2 * (5k - 3) - 2 * (k - 2) Which is the same as... 2 * 5k - 2 * 3 - 2 *k - 2 * 2 or... 10k - 6 - 2k -4 Then you can group all of the like terms. 8k -10 This means that the right side of your equation is equivalent to 8k - 10. So, now you have 5 - (6k + 1) = 8k - 10 The minus sign in front of the set of parenthesis on the left is equivalent to a -1 (in other words, you need to multiply everything in the parenthesis by -1). 5 + -1 * 6k + -1 * 1 = 8k - 10 or 5 - 6k -1 = 8k - 10 Again, you can collect the like terms on the left side yielding: 4 - 6k = 8k - 10 Well, now what do you do? You need to isolate the variable that you are solving for. You can do this by adding 6k to both sides. 4 - 6k + 6k = 8k - 10 + 6k or 4 = 14k - 10 What!? A variable is still next to a constant! Bah! Add 10 to fix this case! 4 + 10 = 14k - 10 + 10 or 14 = 14k Then you just need to divide both sides by 14, yielding: k = 1. Let me know if you have any problems understanding this.

OpenStudy (anonymous):

(5k-3)-(k-2) What we need to do here is get rid of the brackets and bring the "alike" bits together. Now, you have a 5k in the first bracket and a k in the second bracket and the minus sign between means we are going to subtract the second one from the first one so 5k -k = 4k With me so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Now we are going to do the same thing with the numbers. We have a -3 in the first bracket and a -2 in the second so -3 - (-2) is?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

What is -3 -2 ?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

I got up to 0= 10k^2-8k+8. How would you break that down?

OpenStudy (anonymous):

What is 3 -2? and -3 -2 ?

OpenStudy (anonymous):

@Ventricate Why don't you post that as a question?

OpenStudy (anonymous):

3-2 is 1 & -3 -2 is -1

OpenStudy (anonymous):

oh because i thought its related to this question and i got stuck trying to do it.

OpenStudy (anonymous):

OK, some rules 1+2 = 3 1-2 = -1 -1+2 = 1 -1-2 = -3 -1-(-2) = 1 Do you understand the last one?

OpenStudy (anonymous):

Some problem with the last one?

OpenStudy (anonymous):

Or some problem with the others as well?

OpenStudy (anonymous):

no I understand

OpenStudy (anonymous):

OK, let's go back to where we were: Now we are going to do the same thing with the numbers. We have a -3 in the first bracket and a -2 in the second so -3 - (-2) is?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Yes!:-)

OpenStudy (anonymous):

So, to recap, we started with (5k-3)-(k-2) and we found 5k -k = 4k and we found -3 - (-2) = -1 So now we put them together 4k-1 Alright?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now let's rewrite our original question and put in what we found. 5-(6k+1)=2[(5k-3)-(k-2)] becomes 5-(6k+1)=2(4k-1) I changed the [ to ( Doesn't look so bad now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Good, let's repeat the game for what's left. We will take the left hand side (lhs) of the equation 5-(6k+1) is?

OpenStudy (anonymous):

k=-2?

OpenStudy (anonymous):

There's no equals...

OpenStudy (anonymous):

We are doing the same as before, just making it simpler.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5-(6k+1) is?

OpenStudy (anonymous):

Pretend there is a bracket around the 5 if that helps...

OpenStudy (anonymous):

-2k

OpenStudy (anonymous):

There is only one bit that has a k in it, 6k so how did you get 2k?

OpenStudy (anonymous):

Or -2k

OpenStudy (anonymous):

uggggg I hate these probs I added in pareth and then subtracted

OpenStudy (anonymous):

But you just did one that was harder...

OpenStudy (anonymous):

Do you agree that 5-(6k+1) is the same as 5 - 6k -1?

OpenStudy (anonymous):

How about I rewrite it as (0k + 5) - (6k +1) the same way we did it before?

OpenStudy (anonymous):

Remember? We have 0k in the first bracket and 6k in the second bracket and the minus between means we are going to subtract the second from the first etc etc.

OpenStudy (anonymous):

Tired?

OpenStudy (anonymous):

no my comp is being dumb sorry.

OpenStudy (anonymous):

OK, I have a suggestion. http://openstudy.com/groups/mathematics/updates/4e1535780b8b62a9006ea1da#/groups/mathematics/updates/4e189b0b0b8bc2275746e59f is the link for this thread. Why don't you take some time and read through it a couple times until u feel comfortable with it. Then we can continue when u are ready?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok how can I get ahold of you?

OpenStudy (anonymous):

I will find u when u r on (or vice versa) I'm on most days.

OpenStudy (anonymous):

ok thank you for your time.

OpenStudy (anonymous):

If u put a reply in here I will be notified automatically

OpenStudy (anonymous):

Cu soon...ciao!